OpenStudy (anonymous):

Decompose 8y+7/y^2+y-2 into partial fractions.

OpenStudy (merchandize):

Factoring the denominator, we have: y^2 + y - 2 = (y + 2)(y - 1) Then using partial fractions: (8y + 7)/(y^2 + y - 2) = (8y + 7)/[(y + 2)(y - 1)] = A/(y + 2) + B/(y - 1) Multiply both sides by (y + 2)(y - 1) to obtain: 8y + 7 = A(y - 1) + B(y + 2) Expand the right side and collect like terms: 8y + 7 = A(y - 1) + B(y + 2) 8y + 7 = Ay - A + By + 2B 8y + 7 = Ay + By - A + 2B 8y + 7 = (A + B)y + (-A + 2B) Equate the coefficients, so you have the following system of equations: A + B = 8 -A + 2B = 7 The solution is A = 3, B = 5, so we have: (8y + 7)/(y^2 + y - 2) = (8y + 7)/[(y + 2)(y - 1)] = A/(y + 2) + B/(y - 1) = 3/(y + 2) + 5/(y - 1) I hope this helps!